# Bridging the Digital Divide: Small Language Models as a Pathway for Physics and Photonics Education in Underdeveloped Regions

### v1

@@ 1 The Digital Divide in Education in underdeveloped regions
Figure 1: Small Language Models can help bridge the digital divide by enabling interactive on-site science education, bringing AI-powered learning to students without the need for internet access or advanced infrastructure.
@@ 2 The Role and Limitations of Large Language Models (LLMs)
Although LLMs have transformed natural language processing, their real-world deployment faces significant challenges that limit accessibility and sustainability. The computational demands of transformer architectures require expensive GPU clusters for both training and inference, often compelling organizations to depend on cloud-based solutions. While this centralized approach offers convenience, it introduces several critical limitations. A key concern is the escalating energy consumption and infrastructure costs. For instance, pretraining even a moderate-sized model like the 70-billion-parameter Llama 2 emitted 291 tons of CO2 equivalent, which is comparable to the annual emissions of 65 gasoline-powered cars [13, 14]. Scaling to larger architectures exacerbates this problem. For example, training the 405-billion-parameter Llama 3.1 required 30.84 million GPU hours, seventeen times higher than the energy expenditure of Llama 2 (see Figure 3) [15]. Proprietary models such as GPT-4, which are estimated to contain over one trillion parameters, are expected to incur significantly higher environmental costs, although exact figures have not been publicly disclosed by their developers.
Figure 3: CO2 emissions during pretraining. Time: total GPU time required to train each model. Power consumption: peak power capacity per GPU device [13].
@@ 3 Small Language Models (SLMs): Efficiency, Adaptability, and Access
Figure 7: Performance of Qwen2.5-Math models versus general-purpose LLMs on mathematical reasoning benchmarks (GSM8K and MATH) under few-shot chain-of-thought prompting. Specialized smaller models such as Qwen2.5-Math-7B outperform much larger models like LLaMA3.1-405B, while even Qwen2.5-Math-1.5B, with only 1.5 billion parameters, surpasses LLaMA3.1-70B on the MATH benchmark [44].
@@
Figure 9: Comparison of DeepSeek-R1 distilled models and other (non-reasoning) models. Despite their smaller size, the distilled variants of DeepSeek-R1 outperform or match significantly larger models like GPT-4o [60].

### v2

@@ 1 The Digital Divide in Education in underdeveloped regions
Figure 1: Small Language Models can help bridge the digital divide by enabling interactive on-site science education, bringing AI-powered learning to students without the need for internet access or advanced infrastructure. Illustration by S. Cook-Ordonez.
@@ 2 The Role and Limitations of Large Language Models (LLMs)
Although LLMs have transformed natural language processing, their real-world deployment faces significant challenges that limit accessibility and sustainability. The computational demands of transformer architectures require expensive GPU clusters for both training and inference, often compelling organizations to depend on cloud-based solutions. While this centralized approach offers convenience, it introduces several critical limitations. A key concern is the escalating energy consumption and infrastructure costs. For instance, pretraining even a moderate-sized model like the 70-billion-parameter Llama 2 emitted 291 tons of CO2 equivalent, which is comparable to the annual emissions of 65 gasoline-powered cars [13, 14]. Scaling to larger architectures exacerbates this problem. For example, training the 405-billion-parameter Llama 3.1 required 30.84 million GPU hours, seventeen times higher than the energy expenditure of Llama 2 (see Table 1) [15]. Proprietary models such as GPT-4, which are estimated to contain over one trillion parameters, are expected to incur significantly higher environmental costs, although exact figures have not been publicly disclosed by their developers.
Model Size Time (GPU hours) Power (W) Carbon Emitted (tCO2 eq)
7 B 184 320184320184\,320184 320 $400$ $31.22$
13B 368 640368640368\,640368 640 $400$ $62.44$
34B 1 038 33610383361\,038\,3361 038 336 $350$ $153.90$
70B 1 720 32017203201\,720\,3201 720 320 $400$ $291.42$
Total 3311616 539.00
Table 1: CO2 emissions during pretraining for Llama 2 models. Time: total GPU time required to train each model. Power consumption: peak power capacity per GPU device [13].
@@ 3 Small Language Models (SLMs): Efficiency, Adaptability, and Access
Model English Chinese
GSM8K MATH MMLU-STEM CMATH Gaokao Cloze Gaokao QA
8-shot 4-shot 4-shot 6-shot 5-shot 4-shot
General-Purpose Language Models
LLaMA3.1-8B 56.7 20.3 53.1 51.5 8.5 28.5
LLaMA3.1-70B 85.5 41.4 78.1 75.5 11.9 43.3
LLaMA3.1-405B 89.0 53.8
Qwen2-1.5B 58.5 21.7 44.8 55.6 12.7 35.6
Qwen2-7B 79.9 44.2 67.6 76.7 37.3 51.6
Qwen2-72B 89.5 51.1 79.9 85.4 55.9 72.6
Math-Tuned Specialized Models
Qwen2-Math-1.5B 71.3 44.4 50.4 79.6 37.3 50.7
Qwen2-Math-7B 80.4 50.4 65.7 83.2 48.3 57.3
Qwen2-Math-72B 89.1 60.5 79.1 86.4 72.9 69.5
Qwen2.5-Math-1.5B 76.8 49.8 51.3 83.0 47.5 54.1
Qwen2.5-Math-7B 91.6 55.4 67.8 85.0 57.6 69.5
Qwen2.5-Math-72B 90.8 66.8 82.8 89.7 72.9 86.3
Table 2: Performance of Qwen2.5-Math models versus general-purpose LLMs on mathematical reasoning benchmarks (GSM8K and MATH) under few-shot chain-of-thought prompting. Specialized smaller models such as Qwen2.5-Math-7B outperform much larger models like LLaMA3.1-405B, while even Qwen2.5-Math-1.5B, with only 1.5 billion parameters, surpasses LLaMA3.1-70B on the MATH benchmark [44].
@@
Model AIME 2024 MATH-500 GPQA LiveCode CodeForces
Diamond Benchmark
pass@1 cons@64 pass@1 pass@1 pass@1 rating
GPT-4o (0513) 9.3 13.4 74.6 49.9 32.9 759
Claude 3.5 Sonnet (1022) 16.0 26.7 78.3 65.0 38.9 717
Distill-Qwen-1.5B (DeepSeek-R1) 28.9 52.7 83.9 33.8 16.9 954
Distill-Qwen-7B (DeepSeek-R1) 55.5 83.3 92.8 49.1 37.6 1189
Distill-Qwen-14B (DeepSeek-R1) 69.7 80.0 93.9 59.1 53.1 1481
Distill-Qwen-32B (DeepSeek-R1) 72.6 83.3 94.3 62.1 57.2 1691
Distill-LLaMA-8B (DeepSeek-R1) 50.4 80.0 89.1 49.0 39.6 1205
Distill-LLaMA-70B (DeepSeek-R1) 70.0 86.7 94.5 65.2 57.5 1633
Table 3: Comparison of DeepSeek-R1 distilled models and other (non-reasoning) models. Despite their smaller size, the distilled variants of DeepSeek-R1 outperform or match significantly larger models like GPT-4o [60].